The illustration shows a lamp and a candle illuminating a screen positioned between them. If the lamp is moved to a distance r = 1.02 m, how much brighter must the lightbulb be than the candle to have the same illuminance on the screen?
1
Expert's answer
2020-02-07T10:34:02-0500
"\frac{I_2}{I_1}=\frac{r_2^2}{r_1^2}=\frac{1.02^2}{0.34^2}=9"